OpenStudy (anonymous):

A zebra starts from rest and accelerates at 1.9 m/s2. How far has the zebra gone after 5 seconds?

OpenStudy (anonymous):

First, find the final speed: Vf = Vi + at Vi = 0 m/s a = 1.9 m/s^2 t = 5 Vf = 0 + (1.9)(5) = 9.5 m/s Second, find the distance travelled: Vf^2 = Vi^2 + 2a(Xf - Xi) Vf = 9.5 m/s Vi = 0 m/s a = 1.9 m/s^2 Xi = 0 m 9.5^2 = 0 + (2)(1.9)(Xf - 0) 90.25 = 3.8 Xf Xf = 23.75 m The zebra goes 23.75 m.

OpenStudy (amistre64):

s(t) = 1/2 a t^2 + vo t + so since velocity at rest is 0, and the starting position is 0 s(t) = 1/2 a t^2, for some constant acceleration, a